OpenStudy (anonymous):

how do you find all solutions for the equation sin (4x)=1/2

OpenStudy (anonymous):

\[ \sin(4x) =\frac 12 =\sin(\pi/6) \]

OpenStudy (anonymous):

\[ 4 x = \frac \pi 6 + 2 k\pi \\ 4 x = \pi -\frac \pi 6 + 2 k\pi \]

OpenStudy (anonymous):

\[x = \frac \pi {24} + k\frac \pi 2 \\ x =\frac{ 5 \pi} {24} + k\frac \pi 2 \]

OpenStudy (anonymous):

i forgot to put over the interval of [0.2π) so this is 2 of them but not all of them but i think i get how to do the rest thankyou

OpenStudy (anonymous):

yw